# Evaluating and Steering Modality Preferences in Multimodal Large Language Model

### v1

@@ 4 What Modality Preference Has MLLMs Developed?
We investigate the modality preferences of MLLMs by addressing two key questions: Which modality do MLLMs prioritize? and Can modality preference be controlled? We first analyze the inherent modality preferences of different MLLMs, and then delve into how such preferences can be controlled. This investigation helps uncover the underlying mechanisms of modality preference and enables us to apply these insights to downstream tasks.
@@ 4.1 Which Modality Do MLLMs Prioritize?
Before evaluating modality preference, we first assess the ability of MLLMs to answer questions accurately given a single-modality context in the MC2 dataset. All models achieve over 95% accuracy when provided with either textual or visual context. Please refer to Appendix C.1 for detailed results.
Figure 2: Results of modality preference. Left: averaged modality preferences of different MLLMs, Right: vision Ratios for different MLLMs across various tasks.
@@
How do different MLLMs exhibit preferences across various tasks? As described in Section 3.3, we quantify modality preference using model predictions categorized as Vision, Text, and Others, and use the Vision Ratio to quantify the preference for the vision modality, with the results presented in Figure 2 and Table 1. We observe that all MLLMs demonstrate some level of modality preference bias. We observe that the majority of MLLMs exhibit a strong preference for the text modality, particularly as evidenced by the Vision Ratio of LLaVA1.5-13B, which is only 14.4% in the left panel of Figure 2. Interestingly, the Qwen2.5VL and InternVL3 models show a certain degree of preference for the vision modality, which contrasts with the previous conclusion that MLLMs suffer from a severe language prior [23, 37, 45].
A more detailed analysis of different tasks reveals that Qwen2.5VL and InternVL3 have higher Vision Ratios across all tasks compared to other models, especially in high-level recognition tasks such as sports and object recognition, as shown in the right panel of Figure 2. Moreover, we find that low-level tasks (e.g., attribute and color recognition) as well as tasks involving more complex reasoning (such as sentiment analysis, positional reasoning, and counting) consistently exhibit a stronger reliance on the text modality across all MLLMs.
@@
What is the relationship between modality preference and model size? We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference. As shown in the left panel of Figure 3, we observe that for all model families, the preference for the vision modality increases with the model size. And the Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases. However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their size increases.
Figure 3: Evaluation results of modality preference and attention analyse. The left shows modality preference of MLLMs as parametersize changes, measured by Vision Ratio. A higher Vision Ratio indicates a stronger vision modality preference. the right shows the trends of Vision Ratio and multimodal Attention Ratio across different tasks and models.
@@ 4.2 Can Modality Preference Be Controlled?
We employ two intervention methods, content quality manipulation and instruction-based guidance, to explore whether modality preference can be controlled.
Figure 4: Effects of different interventions on the modality preference, measured by Vision Ratio. The left two panels: modifying input quality, including image noise and text grammar errors. The right two panels: using specific instructions to guide MLLMs toward the vision or text modality.
Further analysis using attention attribution. To further investigate the internal mechanisms underlying modality preference, we analyze the attention distributions over the vision and textual contexts in Qwen2.5VL-7B and LLaVA-OneVision-7B. Specifically, we compute the average attention scores for each modality at every token position across all transformer decoder layers of the MLLMs, and define the ratio between visual and textual attention as the Attention Ratio. We then visualize the Attention Ratio alongside the Vision Ratio across a set of representative tasks. As shown in right panel of Figure 3, the trends of Vision Ratio and Attention Ratio are highly aligned across both models and tasks. This suggests that modality preferences are closely related to the internal attention allocation. A similar observation has also been reported in prior work [45].
How does modality content quality affect modality preference? We manipulate the quality of input modality content by introducing Gaussian noise to visual contexts and adding grammatical or spelling errors to textual contexts to observe shifts in modality preference. As shown in the left two panels of Figure 4, we observe that adding noise to images causes a shift in modality preference toward the text modality, whereas introducing grammatical or spelling errors into the text leads to a shift toward the vision modality. It is noteworthy that, despite the degradation in the quality of one modality, the models’ ability to comprehend information from a single modality remains largely unaffected.
This implies that MLLMs may implicitly associate modality reliability with surface-level quality signals (e.g., noise or errors), which then influences their modality weighting during reasoning. Details of the data generation, experimental design, and full results are provided in Appendix C.2.
@@
Figure 5: PCA visualization of modality preference directions under different intervention strategies. Left: representation shifts under noise-based interventions (image noise or textual errors). Middle: representation shifts under instruction-guided interventions. Right: layer-wise absolute difference and standard deviation of hidden states between image-guided and text-guided instruction.
Can modality preference be guided through instruction design? We further investigate the impact of instruction design on modality preference. We design the following task instruction:
USER: <vision context> <textual context> <question>. In this case, there is a conflict between the vision and textual context, and you should rely on the textual/(vision) context rather than the vision/(textual) context.  ASSISTANT: <output>
As shown in the right two panels of Figure 4, instructions directing the model toward either the textual or vision modality effectively guide the modality preference in the specified direction. Moreover, guidance toward the textual modality tends to produce a more pronounced effect compared to guidance toward the vision modality. This tendency may reflect the disproportionate exposure to textual instructions during training, leading to stronger alignment with text-guided cues.
Modality Preference Direction in Representation Space To further understand how the two intervention methods influence modality preference internally, we analyze the hidden representations of the models. Specifically, we apply Principal Component Analysis (PCA) to the hidden states across different layers to identify the dominant direction corresponding to modality preference shifts. Detailed implementation is described in Appendix C.3.
As shown in the left panel of Figure 5, introducing Gaussian noise to visual inputs leads to a noticeable shift in the representation space along a clear direction, while adding grammatical or spelling errors to text inputs results in a comparatively weaker shift. This asymmetry is consistent with our earlier observations in Section 4.2, where textual perturbations steer less pronounced changes in modality preference compared to visual perturbations. Furthermore, the middle panel of Figure 5 illustrates that instruction-based interventions also steer significant representation shifts, aligning with the modality specified in the instruction. Notably, the directionality in the PCA space reveals that the model’s internal states are indeed sensitive to modality control cues, whether they originate from instruction design. Inspired by this, in the following section 5, we propose a probing and Steering method to amplify modality preference and apply it to downstream tasks.
@@
Figure 6: Overall framework of the proposed method. Modality Preference Probing collects the neural activity representations in response to each modality preference prompt. It then computes and scales the direction of modality preference that expresses a preference for a specific modality. Modality Preference Steering selects the target layer during the second inference and adds the scaled modality preference direction to the representation at the corresponding layer at each inference step.
5 Representation-Based Modality Steering
In this section, we adapt the concept of representation engineering [56] to probe the modality preference direction and reveal the consistent representation behavior behind both interventions discussed in Section 4.2. Inspired by this, we steer and amplify the preference direction to control the model’s behavioral expression. Representation engineering involves collecting neural activity by constructing vectors of activation values that cause desired changes to output text, and controlling the model’s behavior by adding the probing vectors to the forward passes of a frozen LLM. As shown in Figure 6, the proposed framework consists of two parts: Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
@@ 5.1 Modality Preference Probing
Modality Preference Probing (MPP) is proposed to probe and collect neural activity that represents the direction of modality preference. Inspired by the pre-training objective of decoder-only MLLMs, Next Token Prediction objective [21] and previous works that extract classification features from the last token [11], we collect neural activity from the last token in the input text.
The process involves probing modality preference through two pairs of modality preference requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context rather than the text context’) and another with a text preference probing (e.g., ‘answer the question based on the text context rather than the vision context’). Let us denote these two inputs by $q^{v}$ (based on the vision context) and $q^{t}$ (based on the text context), and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$. Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states on the two queries at the last token of the input at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM. We identity the direction of modality preference by computing the difference in the hidden states between the paired inputs. More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards expressing the text modality at layer $l$ for a given query as:
@@
Averaging over different queries allows us to capture the activation values most closely associated with modality preference requests, independent of questions.
As shown in the right panel of Figure 5, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples. We observe that layers 20–23 exhibit both higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers. Similar patterns are observed for Qwen2VL and LLaVA-OneVision, as detailed in Appendix D.1. Based on this observation, we select the corresponding layers of the model to control the direction of modality preference in Section 5.2.
Preference Model Sport Attribute Sentiment Positional Counting Color Activity Object Avg
Vision MLLM-only 26.4 12.4 0.8 13.2 4.0 16.0 11.6 38.0 15.3
Inst Design 60.8 24.0 20.0 20.4 10.8 32.0 27.2 63.2 32.3
CoT Prompting 57.6 27.6 16.0 18.0 21.6 35.6 44.4 52.4 34.2
Few shot 32.0 12.0 10.0 11.0 2.0 25.0 9.0 38.0 17.2
ours 78.8 35.6 38.8 29.6 22.4 56.4 45.2 78.4 48.1
Text MLLM-only 12.8 46.0 68.8 38.0 39.6 20.0 43.6 14.0 35.4
Inst Design 14.4 46.8 72.8 40.4 55.6 24.4 35.6 11.6 37.7
CoT Prompting 27.2 63.6 83.6 62.8 75.6 53.2 58.0 20.4 55.6
Few shot 21.0 77.0 89.0 73.0 73.0 60.0 70.0 42.0 63.1
ours 69.6 67.6 84.4 50.8 82.8 57.6 54.8 41.2 63.6
Table 2: Performance by steering Qwen2VL-7B towards vision and Qwen2.5VL-7B towards text.
@@ 5.2 Modality Preference Steering
After probing the direction of the modality preference, we compute the steering vector by re-scaling the direction vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$. The scaling process must carefully balance two objectives: it must be strong enough to effectively steer the model’s modality preference, while simultaneously preserving the model’s normal output behavior. In our preliminary experiments, we observe that setting the weight too large led to repetitive and meaningless outputs, whereas a weight that is too small fails to steer any noticeable change in the model’s modality preference. Unlike previous approaches [56, 38] that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original hidden state distribution. This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities. Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
@@
Finally, during a new round of inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ (as identified in Section 5.1) at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens—including vision tokens, text tokens, and newly generated tokens—to steer the model’s response toward the text modality.
Similarly, to steer responses towards the vision modality, the adjustment is performed in the opposite direction. In the final implementation of our method, no additional data or labels are introduced during either the probing or steering stages. We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the model’s modality preference. Notably, the use of unlabeled or self-generated signals avoids annotation bias, which is a critical property when aiming to enhance the generalization ability of different models across diverse domains.
5.3 Verifying the Effectiveness of Steering Modality Preference
We verify the effectiveness of the proposed method in controlling modality preference on MC2 across several MLLMs, including Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-14B. We consider several training-free approaches as baselines: MLLM-only refers to MLLM directly reasoning in modality-conflicting contexts; Inst-Design uses instructions to guide modality preference direction; CoT-prompting enables complex reasoning through intermediate steps; and Few-shot uses a few examples to guide the model in answering. For detailed method implementation and the results of LLaVA-OneVision-7B and InternVL3-14B, refer to Appendix D.2. As shown in Table 2, the proposed method significantly outperforms the baseline approaches, particularly in high-level recognition tasks.
Phd Model Attribute Sentiment Positional Counting Object Avg
Phd-icc MLLM-only 10.0 2.50 3.50 6.0 8.0 6.0
Inst Design 14.5 2.5 1.5 5.5 25.0 9.8
CoT Prompting 5.0 6.0 8.50 6.5 40.5 13.3
Few shot 3.0 0.5 1.5 5.0 2.0 2.4
+ours 10.0 11.0 14.0 5.0 51.5 18.4
Phd-iac MLLM-only 28.5 8.5 20.5 30.5 50.0 27.6
Inst Design 34.5 13.0 26.0 39.0 60.0 34.5
CoT Prompting 15.5 23.5 30.2 17.0 59.0 29.0
Few shot 17.0 9.0 14.5 29.0 37.0 21.3
+ours 29.5 14.5 35.0 28.5 63.5 34.2
Table 3: Performance of the proposed method on Qwen2VL-7B using the hallucination evaluation dataset from Phd [30]. Phd-icc provides an incorrect context in addition to the correct image, while Phd-iac introduces an inaccurate context to assess model robustness under misleading input.
@@ 5.3 Verifying the Effectiveness of Steering Modality Preference
      +ours 10.22 (+1.3) 19.89 (+1.33)
Table 4: The translation results for Ambigcaps using different instruction for preference towards vision modality and text modality and the proposed method steering modality preference towards text modality. We report the Bleu score for the bidirectional translation between English (en) and Turkish (tr).
5.3.1 Downstream Task Applications of Modality Preference Control
We also evaluate the effects of adjusting modality preference towards specific directions on two downstream tasks. Specifically, we assess the hallucination mitigation effect on two subsets of the PhD dataset [30], namely Phd-icc and Phd-iac. Additionally, we evaluate the performance on a multimodal machine translation (MMT) task using the AmbigCaps [25] dataset.
By steering modality preference towards the vision modality, which places greater trust in visual information, we mitigate multimodal hallucinations. As shown in Table 3, our method significantly outperforms the baseline methods on PhD. Conversely, by steering towards the text modality, the model focuses more on the textual modality, which helps reduce hallucinations in multimodal machine translation. This approach prevents the model from overly focusing on vision information and adding unnecessary objects or other details in the translation output. We use our method to steer the modality preference towards the text modality. As shown in Table 4, proposed method improves the performance of the multimodal machine translation task for the bidirectional translation between English and Turkish. For detailed implementation and results, refer to Appendix D.3.

### v2

@@ 4 Modality Preferences in MLLMs
This section presents a systematic investigation into modality preference in MLLMs, structured around four key research questions: 1) Which modality do MLLMs prioritize? 2) What factors drive these preferences? 3) Can the Vision Ratio provide guidance for downstream task performance? 4) Can modality preference be controlled? This investigation helps uncover the underlying mechanisms of modality preference and enables us to apply these insights to downstream tasks.
Figure 2: Results of modality preference across different MLLMs. Left: Quantified scores for Vision, Others Text modalities using $S_{vision}$, $S_{others}$ and $S_{text}$ as well as Vision Ratio. Right: Trends of Vision Ratio with respect to model parameter size across different MLLMs.
@@ 4.1 Which Modality Do MLLMs Prioritize?
We use the MC2 benchmark to evaluate the modality preferences of 20 open-source MLLMs and the proprietary ChatGPT-4o-mini (Hurst et al., 2024a), detailed in Appendix C.1.
@@
Different MLLMs exhibit different modality preferences. As described in Section 3.3, we quantify modality preference using Vision Ratio, with the results presented in the left panel of Figure 2 and detailed in Table 6. We observe that all MLLMs exhibit clear modality preference, with most models showing a strong preference for text; for instance, LLaVA1.5-7B attains only a 13.4% Vision Ratio. This aligns with the previous findings that MLLMs suffer from a severe language prior (Lee et al., 2024; Parcalabescu & Frank, 2024; Wu et al., 2025). Interestingly, the Qwen2.5VL and InternVL3 show a certain degree of preference towards the vision modality.
@@
Larger MLLMs exhibit stronger preferences for the vision modality. We evaluate models from the LLaVA1.5, LLaVA-Next, Qwen2.5VL, InternVL3, and LLaVA-OneVision families to investigate the relationship between model size and modality preference. As shown in the right panel of Figure 2, we observe that for all model families, the preference for the vision modality increases with the model size. And the Qwen2.5VL and InternVL3 models exhibit a significant preference for the vision modality once the model size increases. However, LLaVA1.5, LLaVA-Next, and LLaVA-OneVision models maintain a noticeable preference for the text modality as their sizes increase.
Vision Ratio aligns with human preference. We further verify whether the Vision Ratio can serve as a human-level measure of modality preference in MLLMs. We randomly sample 100 instances from MC2 and compute the Vision Ratio of four representative models—Qwen2.5VL-7B, LLaVA-OneVision-7B, InternVL3-14B, and LLaVA1.5-7B. In addition, we craft prompts to elicit explicit reasoning chains from the models, specifically targeting their reliance on visual or textual information. To ensure labeling reliability, three expert annotators independently annotate the expressed modality preference for each response, with the final label determined by majority vote. The automatically obtained Vision Ratio scores ([56.3%, 24.6%, 52.3%, 13.9%]) are highly consistent with the ones given by human ([61.0%, 22.0%, 51.0%, 16.0%]), with an average discrepancy of only 2.68%. This indicates that the Vision Ratio can act as a reliable, automated proxy for human assessment of modality preference.
Figure 3: Analysis of modality preferences. Left: Trends of Vision Ratio and multimodal Attention Ratio across different models. Middle: Vision Ratio with Respect to the Proportion of Multimodal Conflict-Context Training Data and different MLLMs. Right: Relationship between visual understanding ability, quantified as the average accuracy across seven widely used benchmarks and the modality preference measured by Vision Ratio.
4.2 What Factors Drive These Preferences?
Given that different MLLMs display varying modality preferences, we examine two primary sources of such differences: the internal attention distribution and the training factors.
Different allocation of attention across modalities. We compute the mean attention scores over all token positions from both modalities and define the ratio of visual attention to total attention as the Attention Ratio. By analyzing Qwen2.5VL-7B and LLaVA-OneVision-7B, we observe that the trends of the Attention Ratio closely align with the Vision Ratio across models in the left panel of Figure 3. This alignment suggests that MLLMs distribute attention unevenly between modalities, which in turn contributes to their divergent modality preferences.
Impact of model scale and training data recipe. Through reviewing the technical reports of the evaluated MLLMs, we find that they all adopt a common architecture comprising a vision encoder, an alignment layer, and an LLM. Thus, we hypothesize that the observed preferences mainly arise from two factors: 1) Exposure to more multimodal contexts, especially with conflicting cases, drives more pronounced shifts in modality preference. 2) Larger LLMs are more capable of shifting their preference during training;
To examine these hypotheses, we construct a training dataset containing vision–text conflict contexts and fine-tune Qwen2.5VL-7B/3B and LLaVA1.5-7B with varying proportions of samples with multimodal conflict contexts, adjusting their preferences toward text or vision. We then measured changes with the Vision Ratio. We optimize MLLMs in the opposite direction of their original preferences and measure changes using the Vision Ratio. As shown in the middle panel of Figure 3, increasing the proportion of multimodal contexts consistently leads to larger preference shifts, supporting Hypothesis 1. This suggests that multimodal inputs maybe create more challenging training conditions, leading to stronger shifts of preference. Furthermore, Qwen2.5VL-7B exhibites greater shifts than Qwen2.5VL-3B under the same conditions, supporting Hypothesis 2. This indicates that larger LLMs demonstrate stronger learning ability and adapt more effectively.
4.3 Can the Vision Ratio provide guidance for downstream task performance?
To demonstrate the correlation between the modality preference and performance for downstream tasks, we evaluate the visual understanding abilities of 10 representative MLLMs. Specifically, we compute the average accuracy across 7 widely used benchmarks, as detailed in the Appendix C.2. We then compare the visual understanding abilities with their modality preference measured by Vision Ratio using MC2, as shown in the right panel of Figure 3. The results reveal a strong positive association between the Vision Ratio and visual understanding ability across the evaluated MLLMs. Specifically, Spearman’s rank correlation (Sedgwick, 2014) reaches $\rho=0.964$, demonstrating that the Vision Ratio provides a highly reliable indicator of visual understanding task performance.
@@ 4.4 Can Modality Preference Be Controlled?
We employ instruction guidance to investigate whether modality preference can be controlled, and conduct a latent space representation analysis to examine the mechanisms, underlying the preference adjustment. Details of the experimental design and results are provided in Appendix C.3.
@@
Figure 4: Analysis of modality preference under instruction-guidance. Left: Adjustment of modality preference using instruction-guided control (Inst-vision vs. Inst-text). Middle: Representation shifts under instruction-guided interventions. Right: layer-wise absolute difference and standard deviation of hidden states between different instruction.
Modality preference can be guided through instruction design. We investigate the impact of instruction design on modality preference, for instance, by explicitly directing the model to rely on a specific modality when answering a question. Specifically, we evaluate the modality preference measured by Vision Ratio for four representative MLLMs including Qwen2.5VL-7B, LLaVA-OneVision-7B, Qwen2VL-7B and InternVL3-14B under the text or vision preference instruction (Inst-text, Inst-vision). As illustrated in the left panel of Figure 4, instructions that steer the model toward a particular modality effectively shape its modality preference.
Modality preference direction in representation space. To further understand how the intervention methods influence modality preference internally, we analyze the hidden representations of the models. Specifically, we apply Principal Component Analysis (PCA; Abdi & Williams, 2010) to the hidden states to identify the dominant direction corresponding to modality preference shifts. The middle panel of Figure 4 shows that instruction-based interventions drive clear shifts in representations, aligning with the modality specified by the instruction. The PCA direction further reveals that the model’s internal states are sensitive to modality control cues, which motivates us to develop representation techniques for adjusting modality preference and to apply these insights to downstream tasks in Section 5.
@@
Figure 5: Overall framework of the proposed method. Modality Preference Probing collects the neural activity, computes and scales the direction of modality preference. Modality Preference Steering selects the target layer during the second inference and adds the scaled modality preference direction to the representation at the corresponding layer at each inference step.
5 Representation Based Modality Preference Steering
Inspired by the representation behavior discussed in Section 4.4, we propose to use the representation engineering (Zou et al., 2023) to steer the modality preference, controlling the model’s behavioral expression. As shown in Figure 5, the proposed framework consists of Modality Preference Probing (§5.1) and Modality Preference Steering (§5.2).
@@ 5.1 Modality Preference Probing
We probe and collect neural activity that represents the direction of modality preference. Inspired by the pre-training Next Token Prediction objective of decoder-only MLLMs (Hurst et al., 2024a) and the method to extract classification features (Feucht et al., 2024), we collect neural activity from the last token in the input text. The process involves probing modality preference through two pairs of requests: one with a vision preference probing (e.g., ‘answer the question based on the vision context’) and another with a text preference probing (e.g., ‘answer the question based on the text context’). Let us denote these two inputs by $q^{v}$ (based on the vision context) and $q^{t}$ (based on the text context), and consider a set of $N$ such pairs ($q^{v}_{i},q^{t}_{i}$), $i\in\{1,\dots,N\}$. Let $\mathbf{x}_{i,\ell}^{v},\mathbf{x}_{i,\ell}^{t}\in\mathbb{R}^{d}$ be the hidden states on the two queries at the last token of the input at layer $\ell\in\{1,\dots,L\}$, where $d$ is the dimension of the chosen MLLM. We identity the direction of modality preference by computing the difference in the hidden states between the paired inputs. More formally, we compute a vector $\mathbf{u}_{\ell}\in\mathbb{R}^{d}$ representing the direction towards the text modality at layer $l$ for a given query as:
@@
Averaging over different queries allows us to capture the activation values most closely associated with modality preference, independent of questions. As shown in the right panel of Figure 4, we compute the absolute values and standard deviations of the modality preference direction $\mathbf{u}^{t}_{\ell}$ across different samples. We observe that layers 20–23 exhibit both higher absolute values and lower variance, indicating that the preference direction is more prominent and stable in these layers. Based on this observation, we select the corresponding layer $\ell^{\prime}$ of the model to control the direction of modality preference in Section 5.2. Similar patterns are observed for Qwen2VL-7B, Qwen2.5VL-7B, LLaVA-OneVision and InternVL3, as detailed in Appendix D.1.
@@ 5.2 Modality Preference Steering
After obtaining the probing direction vector, we compute the steering vector by re-scaling the vector $\mathbf{u}^{t}_{\ell}$ with a weight $w\in\mathbb{R}_{d}$. The scaling process must carefully balance two objectives: 1) it must be strong enough to effectively steer the model’s modality preference, 2) it must preserve the model’s normal output behavior. In our preliminary experiments, we observe that setting the weight too large leads to repetitive and meaningless outputs, whereas a too small weight fails to obtain any noticeable change for modality preference. Unlike previous approaches (Zou et al., 2023; Stolfo et al., 2024) that rely on exhaustive search over a validation set to determine the weight, we propose a principled method that aligns the mean of the probed direction distribution with the mean of the original distribution of hidden states. This strategy ensures that the steering remains effective without disrupting the model’s inherent generation capabilities. Formally, the weight is determined by aligning the mean of the probed direction with the central distribution of the original hidden states and the steering vector is computed by:
@@
Finally, during the second inference, we adjust the original hidden states at the selected steering layer $\ell^{\prime}$ at each decoding step by adding $\mathbf{s}^{t}_{\ell^{\prime}}$ to all tokens to steer the model’s response toward the text modality. Similarly, steering models towards vision modality is performed towards the opposite direction. In the final implementation, no additional data or labels are introduced. We only require two consecutive rounds of inference: the first for probing and the second for steering, effectively controlling the modality preference.
Table 1: Performance for steering Qwen2VL-7B and OneVision-7B towards vision modality and steering Qwen2.5VL-7B and InternVL3-8B towards text modality, measured by $S_{vision}$ and $S_{text}$.
Preference Model MLLM InstDesign CoT FewShot Ours
[…]
OneVision-7B 20.9 31.6 36.7 38.2 42.8
[…]
InternVL3-8B 37.5 52.8 53.1 49.8 57.1
Table 2: Multimodal translation results for Ambigcaps (Li et al., 2021). BLEU scores are reported for English (En) $\leftrightarrow$ Turkish (Tr).
@@
      +Ours 10.22 (+1.30) 19.89 (+1.33)
Table 3: Performance of the proposed method on the visual understanding benchmark, Phd (Liu et al., 2024c). we report the accuracy results on the phd-icc/phd-iac.
Model Attribute Sentiment Positional Counting Object Avg
Qwen2VL-7B 10.0 / 28.5  2.5 / 8.5 3.5 / 20.5 6.0 / 30.5  8.0 / 50.0  6.0 / 27.6
+InstDesign 14.5 / 34.5  2.5 / 13.0 1.5 / 26.0 5.5 / 39.0 25.0 / 60.0  9.8 / 34.5
+CoT  5.0 / 15.5  6.0 / 23.5 8.5 / 30.2 6.5 / 17.0 40.5 / 59.0 13.3 / 29.0
+FewShot  3.0 / 17.0  0.5 / 9.0 1.5 / 14.5 5.0 / 29.0  2.0 / 37.0  2.4 / 21.3
+Ours 10.0 / 34.4 11.0 /16.5 14.0 / 28.3 5.0 / 37.4 51.5 / 64.0 18.4 / 36.1
OneVision-7B 11.5/ 20.5  1.5 / 5.0  1.5 / 16.5  6.5 / 28.5 11.0 / 52.0  6.4 / 24.5
+InstDesign 16.0 / 27.0  5.5 / 12.5  6.0 / 31.5 13.5 / 30.5 34.0 / 61.5 15.0 / 32.6
+CoT 17.3 / 28.4  6.2 / 12.9  7.8 / 33.2 13.8 / 30.9 34.5 / 62.1 15.9 / 33.1
+FewShot 17.0 / 28.0  6.0 / 13.0  7.2 / 32.8 13.9 / 31.0 34.8 / 62.3 16.2 / 33.4
+Ours 19.6 / 30.5  7.8 / 13.5 10.3 / 36.4 15.1 / 29.8 35.6 / 63.5 17.7 / 34.7
5.3 Experiments
We verify the effectiveness of the method in controlling modality preference on MC2 and downstream tasks across Qwen2VL-7B, Qwen2.5VL-7B, and LLaVA-OneVision-7B and InternVL3-8B. We consider widely used training-free approaches as baselines: MLLM refers to employing MLLM to directly reason in modality-conflicting contexts; InstDesign uses instructions to guide modality preference direction; CoT enables complex reasoning through intermediate steps; and FewShot uses four examples to guide the models. For detailed implementation and results, refer to Appendix D.
As shown in Table 2, the proposed method consistently outperforms the baseline approaches on MC2 across both settings, demonstrating its effectiveness in adjusting modality preference. We further assess the effectiveness of the proposed method on two types of downstream tasks: 1) multimodal machine translation (MMT) using AmbigCaps (Li et al., 2021), and 2) visual understanding on PhD (Liu et al., 2024c). The latter includes two subsets—PhD-ica, which contains irrelevant textual context, and PhD-icc, which introduces misleading or incorrect textual information—both of which increase the risk of hallucination. In MMT, the task should primarily ground the source-language text while treating visual information as auxiliary. Accordingly, we adjust modality preference toward the text modality. As shown in Table 2, our method yields an improvement of 1.33 BLEU score over the baselines. By contrast, for PhD, we steer the modality preference toward the vision modality to ground the image. The results in Table 3 demonstrate that our approach achieves substantial improvements across both MLLMs. In particular, when applied to Qwen2VL-7B, our method surpasses the best baseline by an average margin of 6.1 percentage points.